# Bayesian Prediction under Moment Conditioning

### v1

@@ Appendix A Proofs of Main Results
The fixed-point radius
$\rho_{n}=\Big(\tfrac{2\log n}{\lambda_{\min}(H^{*})}\Big)^{1/2}$
balances the exponentially small tail mass against the $O(n^{-1/2})$ lattice discretization error, yielding the sharp curvature-sensitive rate $O(n^{-1/2}\operatorname{polylog}n)$. This scaling ensures that the Gaussian approximation remains uniformly valid while adapting to local curvature, expanding in flat regions, and contracting in highly curved regimes.
A.2 Proof of Theorem 3.4
Decompose
$\mu_{n,m}=\mathbb{E}[\mathcal{L}(X_{1:m}\mid P_{n})\mathbf{1}_{\mathcal{W}_{n}}]+\mathbb{E}[\mathcal{L}(X_{1:m}\mid P_{n})\mathbf{1}_{E_{n}\setminus\mathcal{W}_{n}}].$
On $\mathcal{W}_{n}$, by the hypergeometric coupling and the Gaussian localization,
$\|\mathcal{L}(X_{1:m}\mid P_{n})-P_{n}^{\otimes m}\|_{\mathrm{TV}}\leq\frac{m(m-1)}{2np_{\min}^{*}},\qquad\|P_{n}^{\otimes m}-(P^{*})^{\otimes m}\|_{\mathrm{TV}}\leq m\sqrt{|\mathcal{X}|}\,\rho_{n}\,n^{-1/2}.$
The complement contributes $O(n^{-1})$ by Lemma A.1. Combining these yields
$\|\mu_{n,m}-(P^{*})^{\otimes m}\|_{\mathrm{TV}}=O\left(m\sqrt{\tfrac{\log n}{n\,\lambda_{\min}(H^{*})}}\right).$
@@
If $\min_{x}P(x)>0$, then $\|\mathcal{L}(X_{1:m}\mid P_{n}=P)-P^{\otimes m}\|_{\mathrm{TV}}\leq\frac{m(m-1)}{2n\,p_{\min}}$.
A.3 Proof of Corollary 4.1
By Theorem 3.1, $P_{n}=P^{*}+n^{-1/2}Z+O_{p}(n^{-1})$ with $Z\sim\mathcal{N}(0,(H^{*})^{-1})$ on $\mathsf{T}^{*}$. A second-order expansion gives
@@
where $\Delta_{n}=\sqrt{n}(P_{n}-P^{*})$ and $\xi_{m}$ is Gaussian with covariance $(H^{*})^{-1}$. This is the classical LAN expansion of Le Cam and Yang [37], van der Vaart [54], and contiguity follows by Le Cam’s third lemma.
@@ Appendix B Auxiliary Results and Curvature Computation
Appendix C GEE, and Berk’s Theorem
C.1 Proof of Theorem 6.1 (GMM optimality)
Let $\bar{g}_{n}(\theta)=n^{-1}\sum_{i}g(X_{i},\theta)$ and $A_{\theta}=\partial_{\theta}\mathbb{E}[g(X,\theta)]$. By Theorem 3.1, the conditional type law near $P_{\theta}^{*}$ is Gaussian on $\mathsf{T}_{\theta}^{*}$ with precision $H_{\theta}^{*}$. Parameterize $P=P_{\theta}^{*}+n^{-1/2}V_{\theta}v$ and linearize: $A_{\theta}P=\alpha(\theta)+n^{-1/2}A_{\theta}V_{\theta}v+o(n^{-1/2})$.
Conditioning enforces $A_{\theta}V_{\theta}v=\sqrt{n}\,\bar{g}_{n}(\theta)+o_{p}(1)$. The conditional likelihood $\propto\exp\{-\tfrac{1}{2}v^{\top}H_{\theta}^{*}v\}$ selects the minimal-energy solution, yielding
$\log\mathcal{L}_{\mathrm{cond}}(\theta)=\frac{n}{2}\,\bar{g}_{n}(\theta)^{\top}\Big[(A_{\theta}|_{\mathsf{T}_{\theta}^{*}})H_{\theta}^{*-1}(A_{\theta}|_{\mathsf{T}_{\theta}^{*}})^{\top}\Big]^{-1}\bar{g}_{n}(\theta)+O_{p}(n^{-1/2}).$
C.2 Proof of Theorem 6.2 (GEE robustness)
Let $U_{n}(\theta)=n^{-1}\sum_{i}D_{i}(\theta)^{\top}W_{i}(\theta)\{Y_{i}-\mu_{i}(\theta)\}$ and $\theta_{0}$ solve $\mathbb{E}U_{n}(\theta_{0})=0$. Taylor expansion gives $\|U_{n}(\theta_{0})\|=O_{p}(\sqrt{\log n/n})$ and $H_{n}(\tilde{\theta})\to J(\theta_{0})=\mathbb{E}[D_{i}^{\top}W_{i}D_{i}]$. Hence
$\|\hat{\theta}-\theta_{0}\|=O_{p}\left(\sqrt{\tfrac{\log n}{n\,\lambda_{\min}(J(\theta_{0}))}}\right).$
If $aH_{\theta_{0}}^{*}\preceq J(\theta_{0})\preceq bH_{\theta_{0}}^{*}$ then $\lambda_{\min}(J(\theta_{0}))\geq a\,\lambda_{\min}(H_{\theta_{0}}^{*})$, yielding the curvature rate up to $a^{-1/2}$. Asymptotic normality follows by Lindeberg–Feller.
C.3 Berk’s Theorem via Curvature
Theorem C.1 (Berk via curvature).
Let $\theta_{0}=\arg\min_{\theta}D_{D}(P_{0}\|P_{\theta})$ and $H_{\theta_{0}}^{*}$ the information metric at $\theta_{0}$. Under standard conditions,
@@ C.3 Berk’s Theorem via Curvature
hence $\|\theta-\theta_{0}\|_{2}=O_{p}\big(\sqrt{\tfrac{\log n}{n\,\lambda_{\min}(H_{\theta_{0}}^{*})}}\big)$.
Proof.
By the LAN expansion (Corollary 4.1) centered at $\theta_{0}$, Chernoff bounds on the quadratic form yield the exponential tail; integrating against a positive prior gives the posterior bound. ∎
Corollary C.2 (Predictive misspecification).
For $m=o(\sqrt{n\,\lambda_{\min}(H_{\theta_{0}}^{*})/\log n})$,
$\big\|\mu_{n,m}-(P_{\theta_{0}}^{*})^{\otimes m}\big\|_{\mathrm{TV}}=O\left(m\sqrt{\tfrac{\log n}{n\,\lambda_{\min}(H_{\theta_{0}}^{*})}}\right).$
Remark C.3.
This refines classical posterior concentration under misspecification: flat geometry ($\lambda_{\min}\ll 1$) slows rates, while sharp curvature accelerates them.

### v2

@@ Appendix A Proofs of Main Results
The fixed–point radius
$\rho_{n}=\Bigl(\frac{2\log n}{\lambda_{\min}(H^{*})}\Bigr)^{1/2}$
balances the exponentially small tail mass against the $O(n^{-1/2})$ lattice discretization error, yielding the curvature–sensitive rate $O(n^{-1/2}\operatorname{polylog}n)$. This scaling guarantees uniform Gaussian validity while adapting to local curvature: the window expands in flat regions and contracts in highly curved ones.
A.2 Proof of Theorem 3.4: Predictive Collapse
To bound the total variation distance between the constrained predictive law $\mu_{n,m}$ and its deterministic limit $(P^{*})^{\otimes m}$, decompose the expectation according to the localization window $\mathcal{W}_{n}$:
$\mu_{n,m}=\mathbb{E}\bigl[\mathcal{L}(X_{1:m}\mid P_{n})\mathbf{1}_{\mathcal{W}_{n}}\bigr]+\mathbb{E}\bigl[\mathcal{L}(X_{1:m}\mid P_{n})\mathbf{1}_{E_{n}\setminus\mathcal{W}_{n}}\bigr].$
Here $\mathcal{W}_{n}=\{P\in E_{n}:\|P-P^{*}\|_{2}\leq\rho_{n}n^{-1/2}\}$ with $\rho_{n}^{2}=2\log n/\lambda_{\min}(H^{*})$, and $p_{\min}^{*}=\min_{x}P^{*}(x)>0$.
Inside $\mathcal{W}_{n}$, two finite-sample deviations contribute to the total variation error: (i) the hypergeometric correction between sampling with and without replacement, and (ii) the local Gaussian deviation between $P_{n}$ and $P^{*}$.
The first satisfies
$\|\mathcal{L}(X_{1:m}\mid P_{n})-P_{n}^{\otimes m}\|_{\mathrm{TV}}\leq\frac{m(m-1)}{2n\,p_{\min}^{*}},$
by Lemma A.4. The second obeys
$\|P_{n}^{\otimes m}-(P^{*})^{\otimes m}\|_{\mathrm{TV}}\leq m\sqrt{|\mathcal{X}|}\,\rho_{n}\,n^{-1/2},$
from the Gaussian localization of Theorem 3.1. The complement $E_{n}\setminus\mathcal{W}_{n}$ contributes $O(n^{-1})$ by Lemma A.1. Combining these estimates and substituting $\rho_{n}^{2}=2\log n/\lambda_{\min}(H^{*})$ yields
$\|\mu_{n,m}-(P^{*})^{\otimes m}\|_{\mathrm{TV}}=O\!\left(m\sqrt{\frac{\log n}{n\,\lambda_{\min}(H^{*})}}+\frac{m^{2}}{n\,p_{\min}^{*}}\right),$
which proves Theorem 3.4.
@@
If $\min_{x}P(x)>0$, then for all $m\leq n$,
$\|\mathcal{L}(X_{1:m}\mid P_{n}=P)-P^{\otimes m}\|_{\mathrm{TV}}\leq\frac{m(m-1)}{2n\,p_{\min}},$
uniformly for $P\in\mathcal{P}_{n}$.
Proof of Lemma A.4.
Under sampling without replacement, each pair of draws introduces a dependence of order $1/n$. Bounding the covariance between indicator variables $\mathbf{1}\{X_{i}=x\}$ and $\mathbf{1}\{X_{j}=y\}$ yields $\|\mathcal{L}(X_{1:m}\mid P_{n}=P)-P^{\otimes m}\|_{\mathrm{TV}}\leq m(m-1)/(2n\,p_{\min})$, as in standard finite-population coupling arguments. ∎
A.3 Proof of Corollary 4.1: LAN
By Theorem 3.1, the empirical measure satisfies
$P_{n}=P^{*}+n^{-1/2}Z+O_{p}(n^{-1}),\qquad Z\sim\mathcal{N}\!\bigl(0,(H^{*})^{-1}\bigr)\text{ on }\mathsf{T}^{*}.$
A second-order expansion of the log-density ratio gives
@@
where $\Delta_{n}=\sqrt{n}(P_{n}-P^{*})$ and $\xi_{m}$ is Gaussian with covariance $(H^{*})^{-1}$. This is the classical LAN expansion of Le Cam and Yang [40], van der Vaart [58], showing that $\mu_{n,m}$ and $(P^{*})^{\otimes m}$ are contiguous, and completing the proof of Corollary 4.1.
Remark A.5 (Historical significance).
Corollary 4.1 provides a finite predictive analogue of Le Cam’s local asymptotic normality within the discrete, exchangeable setting of this paper. Where classical LAN theory describes the limiting behavior of parametric likelihoods, this result derives the same quadratic expansion from the geometry of constrained exchangeability. The construction recovers Le Cam’s experiment as the limit of a finite predictive law, linking de Finetti’s representation to the Le Cam–Hájek framework through curvature as the common invariant.
@@ Appendix B Auxiliary Results and Curvature Computation
Appendix C GMM, GEE, and Berk’s Theorem
C.1 Proof of Theorem 5.3: GMM optimality
Let $\bar{g}_{n}(\theta)=n^{-1}\sum_{i}g(X_{i},\theta)$ denote the sample moment and $A_{\theta}=\partial_{\theta}\mathbb{E}[g(X,\theta)]$ the Jacobian of the population moment condition. By Theorem 3.1, the conditional law of empirical types near $P_{\theta}^{*}$ is Gaussian on the tangent space $\mathsf{T}_{\theta}^{*}$ with precision matrix $H_{\theta}^{*}$. Writing $P=P_{\theta}^{*}+n^{-1/2}V_{\theta}v$, with $V_{\theta}$ a basis for $\mathsf{T}_{\theta}^{*}$, gives the local expansion
$A_{\theta}P=\alpha(\theta)+n^{-1/2}A_{\theta}V_{\theta}v+o(n^{-1/2}),$
where $\alpha(\theta)=\mathbb{E}[g(X,\theta)]$. Conditioning on the empirical constraint $\bar{g}_{n}(\theta)=0$ enforces $A_{\theta}V_{\theta}v=\sqrt{n}\,\bar{g}_{n}(\theta)+o_{p}(1)$. Because the conditional density of $v$ is proportional to $\exp\{-\tfrac{1}{2}v^{\top}H_{\theta}^{*}v\}$, the likelihood is maximized at the minimal-energy solution of this constraint. Substituting the implied $v$ yields
$\log\mathcal{L}_{\mathrm{cond}}(\theta)=\frac{n}{2}\,\bar{g}_{n}(\theta)^{\top}\Big[(A_{\theta}|_{\mathsf{T}_{\theta}^{*}})H_{\theta}^{*-1}(A_{\theta}|_{\mathsf{T}_{\theta}^{*}})^{\top}\Big]^{-1}\bar{g}_{n}(\theta)+O_{p}(n^{-1/2}),$
which coincides with the GMM criterion up to a curvature-weighted quadratic form. Hence GMM is the optimal quadratic expansion of the conditional (BETEL) likelihood.
C.2 Proof of Theorem 5.4: GEE robustness
Let
$U_{n}(\theta)=n^{-1}\sum_{i}D_{i}(\theta)^{\top}W_{i}(\theta)\{Y_{i}-\mu_{i}(\theta)\},$
where $D_{i}(\theta)=\partial_{\theta}\mu_{i}(\theta)$ and $W_{i}(\theta)$ are working weights. Let $\theta_{0}$ solve $\mathbb{E}[U_{n}(\theta_{0})]=0$. A Taylor expansion gives $\|U_{n}(\theta_{0})\|=O_{p}(\sqrt{\log n/n})$ and $H_{n}(\tilde{\theta})\to J(\theta_{0})=\mathbb{E}[D_{i}^{\top}W_{i}D_{i}]$. Consequently,
$\|\hat{\theta}-\theta_{0}\|=O_{p}\!\left(\sqrt{\frac{\log n}{n\,\lambda_{\min}\{J(\theta_{0})\}}}\right).$
If $aH_{\theta_{0}}^{*}\preceq J(\theta_{0})\preceq bH_{\theta_{0}}^{*}$ for constants $0<a<b<\infty$, then $\lambda_{\min}\{J(\theta_{0})\}\geq a\,\lambda_{\min}(H_{\theta_{0}}^{*})$, and the same curvature-controlled rate follows up to factor $a^{-1/2}$. Asymptotic normality holds by the Lindeberg–Feller theorem, showing that curvature again governs robustness and rate of convergence.
C.3 Proof of Theorem 6.1 and Corollary 6.2: Berk’s Theorem and Prediction Bound
By the finite-sample LAN expansion of Corollary 4.1, the posterior law of $\Delta_{n}=\sqrt{n}(\theta-\theta_{0})$ satisfies
$\log\frac{d\Pi_{n}}{d\Pi_{0}}(\theta)=\ell_{n}(\theta)-\ell_{n}(\theta_{0})=\Delta_{n}^{\top}\xi_{n}-\tfrac{1}{2}\Delta_{n}^{\top}H_{\theta_{0}}^{*}\Delta_{n}+o_{p}(1),$
where $\xi_{n}\sim\mathcal{N}(0,(H_{\theta_{0}}^{*})^{-1})$ and $H_{\theta_{0}}^{*}$ is the information–geometric Hessian at $\theta_{0}$. Applying a Chernoff bound to this quadratic form yields
@@ Appendix C GMM, GEE, and Berk’s Theorem
for constants $c,C>0$ depending on local geometry and the prior, which proves Theorem 6.1.
The predictive version (Corollary 6.2) follows by substituting this contraction rate into Theorem 3.4, yielding
$\|\mu_{n,m}-(P_{\theta_{0}}^{*})^{\otimes m}\|_{\mathrm{TV}}=O\!\left(m\sqrt{\frac{\log n}{n\,\lambda_{\min}(H_{\theta_{0}}^{*})}}\right),$
for $m=o\!\big(\sqrt{n\,\lambda_{\min}(H_{\theta_{0}}^{*})/\log n}\big)$. This completes the proofs of Theorem 6.1 and Corollary 6.2.
Remark C.1 (Historical note).
Berk’s 1966 theorem marked a turning point: it showed that Bayesian posteriors remain coherent even when the model is false, converging to the pseudo-true parameter rather than to truth itself. The present formulation extends this insight from asymptotic inevitability to finite predictive geometry. It shows that misspecification is not merely stochastic error but geometric distortion: curvature, not contradiction, governs how belief collapses around a useful fiction. In this sense, Berk’s result was not an exception to Bayesian learning but its first cautionary proof that convergence alone does not imply understanding.